OpenStudy (anonymous):

find the equation of the line containing the given pair (-2,-4) and (-6,-5)

OpenStudy (texaschic101):

find the slope by using slope formula: slope(m) = (y2 - y1) / (x2 - x1) (-2,-4) x1 = -2 and y2 = -4 (-6,-5) x2 = -6 nd y2 = -5 now sub slope(m) = (-5 - (-4) / (-6 - (-2) slope(m) = (-5 + 4) / (-6 + 2) slope(m) = -1/-4 slope(m) = 1/4 now we use y = mx + b slope(m) = 1/4 you can use either of your points....I will use (-2,-4) x = -2 and y = -4 now we sub -4 = 1/4(-2) + b -4 = - 1/2 + b -4 + 1/2 = b -8/2 + 1/2 = b -7/2 = b your line is : y = 1/4x - 7/2